# MARS-Sep: Multimodal-Aligned Reinforced Sound Separation

### v1

@@ Abstract
Universal sound separation faces a fundamental misalignment: models optimized for low-level signal metrics often produce semantically contaminated outputs, failing to suppress perceptually salient interference from acoustically similar sources. To bridge this gap, we introduce MARS-Sep, a reinforcement learning framework that reformulates separation as decision making. Instead of simply regressing ground-truth masks, MARS-Sep learns a factorized Beta mask policy that is optimized by a clipped trust-region surrogate with entropy regularization and group-relative advantage normalization. Concretely, we sample masks from a frozen old policy, reconstruct waveforms, and update the current policy using clipped importance ratios—yielding substantially more stable and sample-efficient learning. Multimodal rewards, derived from an audio-text-vision encoder, directly incentivize semantic consistency with query prompts. We further propose a progressive alignment scheme to fine-tune this encoder, boosting its cross-modal discriminability and improving reward faithfulness. Extensive experiments on multiple benchmarks demonstrate consistent gains in Text-, Audio-, and Image-Queried separation, with notable improvements in signal metrics and semantic quality. Our code is available at https://anonymous.4open.science/r/MARS-Sep. Sound separation samples are available at https://mars-sep.github.io/.
@@ 1 Introduction
Sound separation (Liu et al., 2022; Dong et al., 2023; Cheng et al., 2025b; Chen et al., 2022; Mahmud & Marculescu, 2024; Ma et al., 2024; Huang et al., 2025a) is a foundational problem in audio processing with broad impact on downstream tasks such as speech recognition (Shi et al., 2022; Kalda et al., 2024), sound event detection (Turpault et al., 2019; Yin et al., 2025), and acoustic scene analysis (Kim & Chang, 2024; Su et al., 2023). Beyond its standalone importance, separation also serves as a powerful data engine: by decomposing mixtures into isolated sources, it enables large-scale augmentation that improves robustness and generalization (Chiu et al., 2021; Yuan et al., 2022; Cheng et al., 2025a; Manilow et al., 2019). In this work we focus on universal, query-conditioned separation—where the query can be audio, text, or image—and ask how to make the separated output not only signal-clean, but also semantically faithful to the user’s intent.
@@
To bridge this gap, we propose MARS-Sep, a reinforcement learning framework that reformulates mask prediction as stochastic decision-making optimized with multimodal rewards. We cast mask generation as an actor-only trust-region optimization over a factorized Beta policy on time-frequency bins. Training uses a clipped surrogate with entropy regularization and normalized advantages, ensuring stable updates. Instead of focusing on low-level sampling details, our approach leverages multimodal rewards that holistically capture signal fidelity, interference suppression, and perceptual quality across audio, text, and visual queries. To provide reliable reward signals and mitigate reward hacking, we further introduce a progressive alignment strategy that fine-tunes the multimodal encoder to enhance cross-modal discrimination and stabilize policy learning.
@@
  • We demonstrate consistent improvements across SDR/SIR/SAR and SI-SDRi, alongside higher CLAP scores and clearer qualitative separation, confirming both signal-level and semantic gains.
@@ 2.1 Universal and Query-Conditioned Sound Separation
Research on isolating sources from complex mixtures has progressed from domain-specific settings—speech separation (Luo & Mesgarani, 2018; 2019; Zeghidour & Grangier, 2021; Subakan et al., 2021) and music source separation (Luo et al., 2017; Rouard et al., 2023a; Luo & Yu, 2023)—toward Universal Sound Separation (USS) (Kavalerov et al., 2019; Wisdom et al., 2020), which aims to decompose arbitrary mixtures without class constraints. Key enablers include permutation invariant training (PIT) for resolving label permutations (Yu et al., 2017; Postolache et al., 2023) and mixture invariant training (MixIT) for leveraging unlabeled mixtures (Wisdom et al., 2020); large-scale resources such as AudioSet (Gemmeke et al., 2017) further catalyzed progress. Beyond audio-only models, integrating visual context (Majumder et al., 2021; Tan et al., 2023) or using class labels as queries (Chen et al., 2022; 2023) expands separation capability. In parallel, Query-Based Sound Extraction (QBSE) reframes the task as extracting user-specified content while suppressing irrelevant sources. By modality, label-queried systems are simple yet closed-set (Chen et al., 2022; Liu et al., 2024); text-queried approaches such as LASS (Liu et al., 2022) enable open vocabulary but face joint-optimization and generalization hurdles; visual queries exploit images for grounding (Michelsanti et al., 2021; Gao & Grauman, 2021; Ye et al., 2024; Pian et al., 2024); and audio queries use exemplars to target abstract or indescribable sounds (Lee et al., 2019; Chen et al., 2022). Recent attempts unify modalities via cross-attention (Chen et al., 2023) or hybrid encoders (Rouard et al., 2023b), though joint training can limit generalization. Representative systems include CLIPSEP (Dong et al., 2022), which leverages visual data to improve text-queried training, and AudioSep (Liu et al., 2024), which couples CLAP (Wu et al., 2023) with a 14k-hour corpus to achieve strong zero-shot performance. Despite these advances, open-vocabulary robustness, multi-polarity operation (extraction and removal within one framework), and scalable multimodal composition (e.g., “dog barking in this image”) remain challenging; query-mixup style training (Cheng et al., 2025b) offers a promising direction but still leaves room for improved semantic alignment and stability.
2.2 Reinforcement Learning for Large Language Models
@@
Reinforcement learning from human feedback (RLHF) has become a central paradigm for aligning large language models with human preferences. Early work (Ziegler et al., 2019; Ouyang et al., 2022) established the pipeline of learning a reward model from pairwise preference data and then applying policy optimization with Proximal Policy Optimization (Schulman et al., 2017), which significantly improved controllability and reliability of LLM outputs. Subsequent methods sought to reduce the complexity and instability of this process. Direct Preference Optimization (DPO) (Rafailov et al., 2023), for instance, eliminates the explicit reward model and optimizes a preference-based objective directly, providing comparable or superior alignment quality with simpler training. More recently, DeepSeek-R1 (Guo et al., 2025) and related studies (Zheng et al., 2025; Yu et al., 2025) explored alternative optimization strategies such as Group Relative Policy Optimization (GRPO), showing improved stability and reasoning capability compared to PPO-based RLHF.
Extensions of these ideas to multimodal systems have further demonstrated the versatility of reinforcement learning for preference alignment. Vision-R1 (Huang et al., 2025b) and R1-VL (Zhang et al., 2025a) applied reinforcement learning with structured rewards to enhance factual grounding and chain-of-thought reasoning in vision-language models. R1-reward (Zhang et al., 2025b) introduced a refined framework for reward modeling that improves semantic fidelity and mitigates reward hacking, thereby strengthening preference alignment across modalities. Complementary efforts, such as Factually Augmented RLHF (Sun et al., 2024) and RLHF-V (Yu et al., 2024), leveraged fine-grained multimodal feedback to reduce hallucinations and improve semantic consistency, highlighting the broader applicability of RLHF beyond text-only domains.
@@ 3.1.2 OmniSep: Unified Omni-modality Sound Separation with Query-mixup
OmniSep (Cheng et al., 2025b) provides the base separation architecture: a frozen ImageBind (Girdhar et al., 2023) encoder maps audio/image/text inputs to a shared feature space and a Separate-Net (U-Net over STFT magnitudes) predicts masks. Let $Q_{A}$ be the audio query, $Q_{V}$ be the visual query, and $Q_{T}$ be the text query. Training in OmniSep mixes queries across modalities via Query-Mixup
@@ 3.2 MARS-Sep: Reinforcement Learning for Multi-source universal sound separation enhancement
To advance beyond the deterministic optimization paradigm of OmniSep, we introduce a reinforcement learning framework that reformulates sound separation as a stochastic decision-making problem with multimodal rewards (Figure 1).
Specifically, the separator produces a mask prediction that parameterizes the new policy ($\alpha_{new},\beta_{new}$), while a snapshot from the previous step provides the old policy ($\alpha_{old},\beta_{old}$). Masks are sampled from the old policy to ensure stable training, and the separated audio is compared against audio/text/video queries in a shared embedding space using a fine-tuned ImageBind encoder with a multimodal fusion module. The cosine similarity between the separated audio embedding and the fused query embedding yields a scalar reward, which is normalized by a running baseline and group-relative scaling to form advantages. These are combined with the log-probability ratio between old and new policies (i.e., $\text{log}\pi_{\text{old}}$ and $\text{log}\pi_{\text{new}}$) under a clipped surrogate objective, supplemented by entropy regularization for exploration and a KL penalty for stability. The current policy is then snapshotted to serve as the old policy in the next iteration.
@@ 4.1 Experiment Setup
We select three representative baselines for comparison. CLIPSep-NIT (the noise-invariant training version released by the authors) employs CLIP embeddings to guide separation with either visual or textual cues. AudioSep adopts large-scale training with text queries to achieve strong open-domain generalization. OmniSep integrates multiple modalities into a unified separation framework, highlighting the potential of multimodal fusion. These baselines jointly cover vision-guided, text-guided, and multimodal approaches, providing a comprehensive and challenging reference for evaluating the effectiveness and generalizability of our method.
@@ 4.2 Main Results
We provide log-mel spectrograms of a representative sample from the test set of VGGSOUND in Figure 3 for result visualization. Compared with the baseline, MARS-Sep suppresses non-target components more selectively, preserving the target’s harmonic ridges and temporal continuity instead of the blocky dropouts visible in the baseline spectrogram. More samples are shown in Appendix E.4.
5 Conclusion
@@
OmniSep (Cheng et al., 2025b) 6.70±0.66 9.04±0.98 13.61±0.77 4.38±0.48 8.98±0.89
@@
OmniSep (Cheng et al., 2025b) 6.66±0.65 10.00±1.05 13.73±0.76 4.43±0.50 8.79±0.89
@@
OmniSep (Cheng et al., 2025b) 12.37±0.85 17.51±1.16 17.96±0.90 9.18±0.79 5.41±0.98
MARS-Sep (ours) 12.91±0.93 17.61±1.17 18.28±0.93 9.85±0.82 5.18±0.01
@@
OmniSep (Cheng et al., 2025b) 13.03±0.96 18.97±1.16 17.88±1.00 10.21±0.89 6.53±1.03

### v2

@@ Abstract
Universal sound separation faces a fundamental misalignment: models optimized for low-level signal metrics often produce semantically contaminated outputs, failing to suppress perceptually salient interference from acoustically similar sources. We introduce a preference alignment perspective, analogous to aligning LLMs with human intent. To address this, we introduce MARS-Sep, a reinforcement learning framework that reformulates separation as decision making. Instead of simply regressing ground-truth masks, MARS-Sep learns a factorized Beta mask policy that is steered by a preference reward model and optimized by a stable, clipped trust-region surrogate. The reward, derived from a progressively-aligned audio-text-vision encoder, directly incentivizes semantic consistency with query prompts. Extensive experiments on multiple benchmarks demonstrate consistent gains in Text-, Audio-, and Image-Queried separation, with notable improvements in signal metrics and semantic quality. Our code is available at https://github.com/mars-sep/MARS-Sep. Sound separation samples are available at https://mars-sep.github.io/.
@@ 1 Introduction
Sound separation (Liu et al., 2022; Dong et al., 2023; Cheng et al., 2025d; Chen et al., 2022; Mahmud and Marculescu, 2024; Ma et al., 2024; Huang et al., 2025b) is a foundational problem in audio processing with broad impact on downstream tasks such as speech recognition (Shi et al., 2022; Cheng et al., 2023; Kalda et al., 2024; Fu et al., 2024), spoken dialogue (Ji et al., 2024; Cheng et al., 2025b; Fu et al., 2025; Cheng et al., 2025a) sound event detection (Turpault et al., 2019; Yin et al., 2025), and acoustic scene analysis (Kim and Chang, 2024; Su et al., 2023). Beyond its standalone importance, separation also serves as a powerful data engine: by decomposing mixtures into isolated sources, it enables large-scale augmentation that improves robustness and generalization (Chiu et al., 2021; Yuan et al., 2022; Cheng et al., 2025c; Manilow et al., 2019). In this work, we focus on universal, query-conditioned separation—where the query can be audio, text, or image—and ask how to make the separated output not only signal-clean, but also semantically faithful to the user’s intent.
@@
Inspired by RLHF which learns a preference model from human data and uses it as a reward signal to steer the base LLM, we conceptualize query-conditioned sound separation as an identical preference alignment problem: the user query (audio, text, or image) is the preference, and the goal is to produce an output that maximizes semantic alignment with the query.
Based on this logic, we reformulate sound separation as a meta-reasoning task. We treat the base separation architecture as a base policy and use reinforcement learning as the optimization algorithm. The human preference is captured by a learned multimodal reward model that provides a high-level semantic signal.
To this end, we propose MARS-Sep, a reinforcement learning framework that reformulates mask prediction as stochastic decision-making optimized with multimodal rewards. We cast mask generation as an actor-only trust-region optimization over a factorized Beta policy on time-frequency bins. Training uses a clipped surrogate with entropy regularization and normalized advantages, ensuring stable updates. Instead of focusing on low-level sampling details, our approach leverages multimodal rewards that holistically capture signal fidelity, interference suppression, and perceptual quality across audio, text, and visual queries. To provide reliable reward signals and mitigate reward hacking, we further introduce a progressive alignment strategy that fine-tunes the multimodal encoder to enhance cross-modal discrimination and stabilize policy learning.
@@
  • We demonstrate consistent improvements across SDR/SIR/SAR and SI-SDRi, alongside higher CLAP scores and clearer qualitative separation for both synthesized and in-the-wild samples, confirming both signal-level and semantic gains.
@@ 2.1 Universal and Query-Conditioned Sound Separation
Research on isolating sources from complex mixtures has progressed from domain-specific settings—speech separation (Luo and Mesgarani, 2018; 2019; Zeghidour and Grangier, 2021; Subakan et al., 2021; Linhui et al., 2023) and music source separation (Luo et al., 2017; Rouard et al., 2023b; Luo and Yu, 2023)—toward Universal Sound Separation (USS) (Kavalerov et al., 2019; Wisdom et al., 2020), which aims to decompose arbitrary mixtures without class constraints. Key enablers include permutation invariant training (PIT) for resolving label permutations (Yu et al., 2017; Postolache et al., 2023) and mixture invariant training (MixIT) for leveraging unlabeled mixtures (Wisdom et al., 2020); large-scale resources such as AudioSet (Gemmeke et al., 2017) further catalyzed progress. Beyond audio-only models, integrating visual context (Majumder et al., 2021; Tan et al., 2023) or using class labels as queries (Chen et al., 2022; 2023) expands separation capability. In parallel, Query-Based Sound Extraction (QBSE) reframes the task as extracting user-specified content while suppressing irrelevant sources. By modality, label-queried systems are simple yet closed-set (Chen et al., 2022; Liu et al., 2024); text-queried approaches such as LASS-Net (Liu et al., 2022) enable open vocabulary but face joint-optimization and generalization hurdles; visual queries exploit images for grounding (Michelsanti et al., 2021; Gao and Grauman, 2021; Ye et al., 2024; Pian et al., 2024); and audio queries use exemplars to target abstract or indescribable sounds (Lee et al., 2019; Chen et al., 2022). Recent attempts unify modalities via cross-attention (Chen et al., 2023) or hybrid encoders (Rouard et al., 2023a), though joint training can limit generalization. Representative systems include CLIPSEP (Dong et al., 2022), which leverages visual data to improve text-queried training, and AudioSep (Liu et al., 2024), which couples CLAP (Wu et al., 2023) with a 14k-hour corpus to achieve strong zero-shot performance. Despite these advances, open-vocabulary robustness, multi-polarity operation (extraction and removal within one framework), and scalable multimodal composition (e.g., “dog barking in this image”) remain challenging; query-mixup style training (Cheng et al., 2025d) offers a promising direction but still leaves room for improved semantic alignment and stability.
Beyond these discriminative and cross-modal systems, several recently proposed models extend query-conditioned separation into generative or flow-based paradigms. More recent generative approaches such as FlowSep (Yuan et al., 2025) employ rectified-flow matching to synthesize query-consistent sources directly in the latent space, and ZeroSep (Huang et al., 2025b) performs zero-training, text-conditioned separation using a pre-trained audio-language diffusion backbone. In contrast to fully synthesis-based models, DAVIS (Huang et al., 2024) and its successor DAVIS-Flow (Huang et al., 2025a) adopts a video-conditioned rectified-flow trajectory that remains time-aligned with the mixture, enabling evaluation under standard signal-level metrics. These developments collectively illustrate a shift toward richer multimodal conditioning and generative reasoning in sound separation, further motivating methods that unify signal fidelity with semantic alignment.
2.2 Reinforcement Learning for Large Language Models and Multimodal Learning
@@
Reinforcement learning from human feedback (RLHF) has become a central paradigm for aligning large language models with human preferences. Early work (Ziegler et al., 2019; Ouyang et al., 2022) established the practice of training a reward model from pairwise preferences and optimizing the policy via Proximal Policy Optimization (PPO) (Schulman et al., 2017). Later methods reduced this pipeline’s complexity, including Direct Preference Optimization (DPO) (Rafailov et al., 2023), which removes the explicit reward model, and more recent Group Relative Policy Optimization (GRPO) approaches (Guo et al., 2025; Zheng et al., 2025; Yu et al., 2025) that improve stability and reasoning over PPO-based RLHF.
These alignment strategies have also extended to multimodal models. Vision-R1 (Huang et al., 2025c) and R1-VL (Zhang et al., 2025a) apply structured rewards to improve grounding and chain-of-thought reasoning, while R1-reward (Zhang et al., 2025b) strengthens multimodal reward modeling to enhance semantic fidelity and robustness. Together, these works highlight the applicability of RLHF-style preference alignment beyond text-only LLMs.
@@ 3.1.2 OmniSep: Unified Omni-modality Sound Separation with Query-mixup
OmniSep (Cheng et al., 2025d) provides the base separation architecture: a frozen ImageBind (Girdhar et al., 2023) encoder maps audio/image/text inputs to a shared feature space and a Separate-Net (U-Net over STFT magnitudes) predicts masks. Let $Q_{A}$ be the audio query, $Q_{V}$ be the visual query, and $Q_{T}$ be the text query. Training in OmniSep mixes queries across modalities via Query-Mixup
@@ 3.2 MARS-Sep: Reinforcement Learning for Multi-source universal sound separation enhancement
As established in our introduction, we frame sound separation as a preference alignment problem, analogous to the RLHF pipeline for LLMs. To operationalize this framework (see Figure 1), we must define three core components:
Base Policy ($\pi_{\theta}$).
The sound separation model that takes the state (mixture spectrogram $X$ and query $Q$) and produces an action (the mask $M$).
Preference Reward Model ($R$).
The model that scores the alignment. We use our progressively-aligned multimodal encoder to provide this scalar reward, measuring semantic consistency between the separated audio and the query.
Optimization Process.
The policy updating algorithm that steers the base policy to maximize the reward. We adapt a stable, trust-region policy optimization algorithm for this purpose.
During the training stage, the separator produces a mask prediction that parameterizes the new policy ($\alpha_{new},\beta_{new}$), while a snapshot from the previous step provides the old policy ($\alpha_{old},\beta_{old}$). Masks are sampled from the old policy to ensure stable training, and the separated audio is compared against audio/text/video queries in a shared embedding space using a fine-tuned ImageBind encoder with a multimodal fusion module. The cosine similarity between the separated audio embedding and the fused query embedding yields a scalar reward, which is normalized by a running baseline and group-relative scaling to form advantages. These are combined with the log-probability ratio between old and new policies (i.e., $\text{log}\pi_{\text{old}}$ and $\text{log}\pi_{\text{new}}$) under a clipped surrogate objective, supplemented by entropy regularization for exploration and a KL penalty for stability. The current policy is then snapshotted to serve as the old policy in the next iteration.
@@ 4.1 Experiment Setup
We select five representative baselines for comparison. CLIPSep-NIT (Dong et al., 2023) (the noise-invariant training version released by the authors) employs CLIP embeddings to guide separation with either visual or textual cues. AudioSep (Liu et al., 2024) adopts large-scale training with text queries to achieve strong open-domain generalization. OmniSep (Cheng et al., 2025d) integrates multiple modalities into a unified separation framework, highlighting the potential of multimodal fusion. LASS-Net (Liu et al., 2022) addresses the language-queried audio source separation task via a joint Transformer-based query encoder and ResUNet separation network, conditioned on natural-language descriptions. Second, iQuery (Chen et al., 2023) formulates instruments as learnable audio-query prototypes and leverages visually-named cross-modal attention to disentangle and separate instrument sounds from videos.
In addition to the above, we further include two generative approach and one zero-shot approach for comparison. FlowSep (Yuan et al., 2025) introduces a generative rectified flow matching (RFM) model in the VAE latent space to synthesize separated audio from noise under language query guidance, thereby enabling cleaner outputs in highly overlapping scenarios. ZeroSep (Huang et al., 2025b) proposes a zero-training audio separation framework that repurposes pre-trained text-guided audio diffusion models to perform open-set language-queried separation without task-specific fine-tuning. We also include DAVIS-Flow (Huang et al., 2025a), a visually-guided generative audio-visual separation framework that applies flow-matching to synthesize separated spectrograms from noise.
@@ 4.2 Main Results
We provide log-mel spectrograms of a representative sample from the test set of VGGSOUND in Figure 3 for result visualization. Compared with the baseline, MARS-Sep suppresses non-target components more selectively, preserving the target’s harmonic ridges and temporal continuity instead of the blocky dropouts visible in the baseline spectrogram.
We additionally report CLAP scores for the generative separation frameworks FlowSep and ZeroSep, which do not produce mask-based or waveform-aligned outputs and are therefore not directly comparable under signal-level metrics. As shown in Table 3, these generative models exhibit substantially higher variance and often inflated CLAP similarity due to their synthesis-based decoding. In contrast, MarsSep provides far more stable and consistent semantic alignment, while also enabling evaluation under standard separation metrics. To further assess the quality of the separated audio, we introduce CLAP audio scores, which measure the semantic similarity between the separated audio and the target audio. This additional evaluation highlights MarsSep’s ability to maintain strong semantic fidelity to the target, even when compared to models with synthesis-based outputs, where CLAP scores tend to be more volatile.
As for ablation studies, the impact of hyperparameter settings, modality query embedding fusion module, progressive finetuning strategy, discrimination ability of multimodal encoder with or without finetuning and other significant factors are discussed in Appendix E.1.
More samples including synthesized audio and in-the-wild audio are shown in Appendix E.6.
@@
DAVIS-Flow (Huang et al., 2025a) 6.60±1.02 8.99±0.93 13.48±0.85 4.32±1.03 6.57±0.94
OmniSep (Cheng et al., 2025d) 6.70±0.66 9.04±0.98 13.61±0.77 4.38±0.48 8.98±0.89
@@
iQuery (Chen et al., 2023) 6.20±0.78 9.59±0.88 13.45±1.01 3.77±0.46 6.08±1.12
DAVIS-Flow (Huang et al., 2025a) 6.52±1.01 9.87±0.98 13.54±0.93 4.32±0.96 8.89±1.02
OmniSep (Cheng et al., 2025d) 6.66±0.65 10.00±1.05 13.73±0.76 4.43±0.50 8.79±0.89
@@
DAVIS-Flow (Huang et al., 2025a) 8.97±0.85 15.99±1.55 17.88±0.95 9.23±1.03 5.53±0.75
OmniSep (Cheng et al., 2025d) 12.37±0.85 17.51±1.16 17.96±0.90 9.18±0.79 5.41±0.98
MARS-Sep (ours) 12.91±0.93 17.61±1.17 18.28±0.93 9.85±0.82 6.18±0.93
@@
i-Query ((Chen et al., 2023)) 12.52±0.84 17.05±1.01 17.65±1.23 9.44±0.99 5.50±1.22
DAVIS-Flow (Huang et al., 2025a) 13.52±1.14 19.00±1.02 17.54±0.93 10.32±0.85 5.99±1.00
OmniSep (Cheng et al., 2025d) 13.03±0.96 18.97±1.16 17.88±1.00 10.21±0.89 6.53±1.03
@@
Table 3: CLAP score comparison (Text-queried) across generative separation frameworks on the MUSIC-clean+ and VGGSOUND-clean+ datasets.
Method Dataset $CLAP_{t}$ score (%) $CLAP_{a}$ score (%)
ZeroSep MUSIC-clean+ $0.02\pm 15.14 22$$2.86\pm 18.55$
FlowSep $10.67\pm 14.17$ $39.25\pm 29.86$
MarsSep (Ours) $6.18\pm 0.93$ $21.56\pm 1.08$
ZeroSep VGGSOUND-clean+ $15.91\pm 14.17$ $22.65\pm 19.98$
FlowSep $8.84\pm 13.27$ $56.07\pm 19.57$
MarsSep (Ours) $9.03\pm 0.94$ $18.70\pm 1.23$
5 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant No. U25B2064, the “Pioneer” and “Leading Goose” R&D Program of Zhejiang under (Grant No. 2025C02110), Public Welfare Research Program of Ningbo under (Grant No. 2024S062), and Yongjiang Talent Project of Ningbo under (Grant No. 2024A-161-G).